How many milliliters of alcohol must be added to 200 milliliters of a 25% iodine solution to make a 10% iodine solution
Answer by ankor@dixie-net.com(22740)

How many milliliters of alcohol must be added to 200 milliliters of a 25% iodine solution to make a 10% iodine solution:
:
Let x = amt of alcohol required;
:
Think of Alcohol is a 0% iodine solution
:
Final mixture amt = (x+200)
:
Write a a % iodine equation:
.25(200) + 0(x) = .10(x+200)
:
50 + 0 = .10x + 20
:
50 - 20 = .10x
:
30 = .10x
:
x = 30/.10
:
x = 300 ml of alcohol required to make a 10% iodine solution
;
:
Check solution:
the total amt 200 + 300 = 500
.25(200) = .1(500)
50 = 50